# STG-Mamba: Spatial-Temporal Graph Learning via Selective State Space Model

### v1

@@ 3.3 GS3B: Graph Selective State Space Blocks
Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1. After being processed by Layer Normalization, the input is divided into two sub-branches. In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning. After that, the learned features are normalized by the Layer Norm. For the second branch, the input STG embedding undergoes a Linear Layer and an activation function. Subsequently, the two branches are fused together by an element-wise dot product operation. Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output. In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function.
@@ 3.4 Computational Efficiency of STG-Mamba
Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods. STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost. In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM. In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM. In common practices we often have $BL^{2}d^{\prime}\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training. Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase. SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 32 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 300. The number of Encoder/Decoder Layer is set as $N=8$ and $N^{\prime}=4$, respectively. We use the historical ground-truth data from the past 12 steps to forecast one future time step. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
@@ 4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2. Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow). In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba. It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods.
@@ 4.3.2 Computational Efficiency Evaluation
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantitively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.

### v2

@@ 3.3 GS3B: Graph Selective State Space Blocks
Figure 2: The selection mechanism of STG-Mamba that facilitates input-dependent feature selection and focusing.
Inspired from Mamba, the GS3B is the major component of STG-Mamba, as depicted in Figure 1. After being processed by Layer Normalization, the input is divided into two sub-branches. In the first branch, the input STG embedding passed through a Linear Layer, followed by the proposed KFGN network and an activation function, before entering the GSS-Layer for selective feature learning. After that, the learned features are normalized by the Layer Norm. For the second branch, the input STG embedding undergoes a Linear Layer and an activation function. Subsequently, the two branches are fused together by an element-wise dot product operation. Finally, the merged embedding passes through a linear layer, and the outcome is combined with a residual connection to form the GS3B block’s output. In this research, Sigmoid Linear Unit (SiLU) is employed as the default activation function. We further comprehend the detailed process of the STG-Mamba’s Selection Mechanism in Figure 2.
@@ 3.4 Computational Efficiency of STG-Mamba
Leveraging a data-specific selection process, the increasing of sequence length $L$ or graph network node scale $N$ by $k$-fold can results in huge expansion of model parameters for Transformer-based methods. STG-Mamba introduces a sophisticated algorithm that is conscious of hardware capabilities, exploiting the layered structure of GPU memory to minimize the ensuing cost. In detail, for a batch size of $B$, STG feature dimension of $d^{\prime}$, and network node number of $N$, STG-Mamba reads the $O(BLd^{\prime}+Nd^{\prime})$ input data of $\mathbb{A},\mathbb{B},\mathbb{C},\Delta$ from High Bandwidth Memory (HBM), processes the intermediate stages sized $O(BLd^{\prime}N)$ in Static Random-Access Memory (SRAM), and dispatches the ultimate output, sized $O(BLd^{\prime})$, back to HBM. In contrast, if a Transformer-based model is employed, the HBM will read $O(BL^{2}d^{\prime})$ input data, and the computational size remains $O(BL^{2}d^{\prime})$ in SRAM. In common practices we often have $BL^{2}d\gg BLd^{\prime}+Nd^{\prime}$, which means SSSM-based approach can effectively decrease the computational overhead in large-scale parallel processing, such as neural network training. Additionally, by not retaining intermediate states, it also diminishes memory usage, since these states are recalculated during the gradient assessment in reverse phase. SSSM’s GPU-optimized design ensures STG-Mamba operates with a linear time complexity $O(L)$ relative to the length of the input sequence, marking a substantial speed advantage over traditional transformers’ dense attention computations, which exhibit quadratic time complexity $O(L^{2})$.
@@ 4.2 Implementation Settings
We split all the datasets with a ratio of 6:2:2 for Training/Validation/Testing sets along the time axis, respectively. Before starting model training session, all the data samples are normalized into range $[0,1]$ with MinMax normalization method. We set the batch size as 32 and employ AdamW as the Optimizer. The learning rate is initialized as $1e^{-3}$, with a weight decay of $5e^{-2}$. CosineAnnealingLR Loshchilov and Hutter (2016) is adopted as the learning rate scheduler, with a maximum of 50 iterations, and a minimum learning rate $1e^{-5}$. Model training epochs is set as 300. The number of Encoder/Decoder Layer is set as $N=4$ and $N^{\prime}=4$, respectively. We use the historical ground-truth data from the past 12 steps to forecast one future time step. Mean Squared Error (MSE) is employed as the Loss Function for optimization. And finally, we employ RMSE, MAE, MAPE as the evaluation metrics. All experiments are conducted under one NVIDIA A100 80GB GPU for neural computing acceleration.
@@ 4.3.1 Experimental Results Comparison on Three STG Forecasting Tasks
A comprehensive evaluation between STG-Mamba and the proposed baseline methods is conducted, and we report the RMSE/MAE/MAPE performance results in Table 2. Our STG-Mamba consistently outperforms other benchmark methods on the three tasks, except for the MAPE criteria of PeMS04 (Flow). In PeMS04 (Flow) forecasting task, STAEformer shows slightly better MAPE results than STG-Mamba. It can also be easily observed that Transformer-based methods which integrate GNNs inside the architecture always gain superior performance than sole GNN-based methods. For instance, PDFormer, STAEformer, and MultiSPANS always show better performance than STGCN, STSGCN, and STG-NCDE. The reason is that although GNN is an effective approach of modeling STG tasks, Transformer-based methods have stronger ability to capture local and long-term dynamic dependencies. Within the GNN-based methods, STG-NCDE and DDGCRN stand out for their competitive performance. STG-NCDE models continuous-time dynamics within STG networks through neural controlled differential equations (NCDE), offering a more precise representation of spatiotemporal feature changes over time compared to other discrete GNN-based methods. Figure 3 presents a direct performance visualization comparison among STG-Mamba and other superior Transformer-based baselines (ASTGNN/PDFormer/STAEformer) on PeMS04, HZMetro, and KnowAir dataset, respectively. We visualize the predictions over 72 consecutive time steps using the average results of multiple 12-step forecast settings. For each dataset, four node individuals are randomly selected (i.e., 4 random traffic sensors of PeMS04, 4 random metro stations of HZMetro, and 4 random weather stations of KnowAir). The results indicate that STG-Mamba consistently yields the prediction curve closest to the ground truth waveform, featuring more accurate STG sequence peak values and fewer fluctuations.
Figure 3: STG forecasting results visualization and comparison among STG-Mamba and other superior Transformer-based baselines.
@@ 4.3.2 Computational Efficiency Evaluation
STG-Mamba offers substantial improvement in computational efficiency in addition to performance gain. In order to quantatively evaluate computational efficiency, we select the widely recognized Floating Point Operations (FLOPs) and Inference Time as criteria. For both of the two criteria, a smaller value indicates better model performance. STG-Mamba is compare with other Transformer-based benchmark methods: ASTGNN, PDFormer, and STAEformer, on the same testing set.
@@
4.4 Robustness Analysis
Table 4: Robustness analysis on PeMS04. We compare the forecasting performance among STG-Mamba and 4 superior baseline methods during rush/non rush hours, weekend/non weekend settings.
Models Rush Hours Non Rush Hours Weekend Non Weekend
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
ASTGNN 33.57 21.05 13.81 29.76 18.22 12.49 30.66 19.08 13.35 31.82 19.53 12.28
PDFormer 31.69 19.54 12.38 28.95 18.16 11.97 29.87 18.26 12.04 30.39 18.68 12.23
STAEformer 31.36 19.10 12.22 29.47 17.98 11.93 29.83 18.19 11.96 30.26 18.42 12.27
MultiSPAN 30.95 19.28 13.31 30.04 18.37 12.54 30.13 18.42 12.73 30.52 19.31 13.34
STG-Mamba 29.58 18.13 12.15 29.23 18.04 11.91 29.31 18.05 11.92 29.57 18.15 12.16
STG data exhibit notable periodicity and diversity. For urban mobility/traffic data, there are distinct differences in traffic characteristics between the peak travel periods in the morning and evening and the none-peak periods. Moreover, the traffic patterns over the 24 hours of weekdays and those of weekends also differ substantially. In light of the variations in STG characteristics due to external environmental changes, it is imperative to assess whether deep learning models can accurately and effectively model the spatial-temporal dependencies under various external conditions.
As such, we establish four distinct external scenarios: (a) Traffic rush hours: Monday to Friday, 8:00-11:00 AM and 16:00-19:00 PM; (b) Non-rush hours: encompassing the remaining time periods from Monday to Friday; (c) Weekend: including all hours on Saturday and Sunday; and (d) Non-weekend: including all time periods from Monday to Friday. Extensive experiments were conducted on the PeMS04 dataset. Table 4 presents the STG forecasting results corresponding to the four traffic scenarios. We compare the prediction performance of STG-Mamba with the other four superior baselines, and report the RMSE/MAE/MAPE metrics.
An ideal deep learning-based STG system should maintain robustness to external disturbances (changes of external environment or events). This is specifically reflected in the fact that when facing different scenarios, the model performance should maintain as little difference as possible, rather than being good at some scenarios but being bad at other occasions.
Among the compared methods, ASTGNN, PDFormer, and STAEformer exhibit relatively large differences in the prediction results between the Rush hour and Non-rush Hour scenarios. In the Weekend/Non Weekend scenario, although the performance differences of each model are not as significant as in the Rush hour/Non-Rush hour scenario, we observe that STG-Mamba still shows the smallest performance difference (the maximum robustness) and achieves the best RMSE/MAE/MAPE metrics. In summary, STG-Mamba proves the best robustness and outperforms existing baseline methods even in face of complex and diversity traffic scenes.